OpenStudy (anonymous):

y varies inversely as twice x. When x = 2, y = 3. Find y when x = 6

OpenStudy (anonymous):

Given: y = k/(2x) x = 2 y = 3 Plug in and solve for k: (3) = k/(2(2)) 3 = k/4 k = 12 Rewrite original equation: y = 12/(2x) Plug in x = 6: y = 12/(2(6)) = 12/12 = 1

OpenStudy (anonymous):

I dot under stand what the answer is 1 2 4 or 6

OpenStudy (anonymous):

1

OpenStudy (anonymous):

oh thanks

OpenStudy (anonymous):

You're welcome.